# DAPO: An Open-Source LLM Reinforcement Learning System at Scale

### v1

@@ Abstract
Figure 1: AIME 2024 scores of DAPO on the Qwen2.5-32B base model, outperforming the previous SoTA DeepSeek-R1-Zero-Qwen-32B using 50% training steps.
@@ 3.1 Raise the Ceiling: Clip-Higher
We propose the Clip-Higher strategy to address this issue. Clipping over the importance sampling ratio is introduced in Clipped Proximal Policy Optimization (PPO-Clip) [21] to restrict the trust region and enhance the stability of RL. We identify that the upper clip can restrict the exploration of the policy. In this case, it is much easier to make an ‘exploitation token’ more probable, than to uplift the probability of an unlikely ‘exploration token’.
Concretely, when $\varepsilon=0.2$ (the default value of most algorithms), consider two actions with probabilities $\pi_{\theta_{\text{old}}}(o_{i}\mid q)=0.01$ and $0.9$. The maximum possible updated probabilities $\pi_{\theta}(o_{i}\mid q)$ are $0.012$ and $1.08$, respectively. This implies that for tokens with a higher probability (e.g., 0.9) is less constrained. Conversely, for low-probability tokens, achieving a non-trivial increase in probability is considerably more challenging. Empirically, we also observe that the maximum probability of clipped tokens is approximately $\pi_{\theta}(o_{i}\mid q)<0.2$ (Figure 3(a)). This finding supports our analysis that the upper clipping threshold indeed restricts the probability increase of low-probability tokens, thereby potentially constraining the diversity of the system.
@@
We increase the value of $\varepsilon_{\text{high}}$ to leave more room for the increase of low-probability tokens. As shown in Figure 2, this adjustment effectively enhances the policy’s entropy and facilitates the generation of more diverse samples. We opt to keep $\varepsilon_{\text{low}}$ relatively small, because increasing it will suppress the probability of these tokens to 00, resulting in the collapse of the sampling space.
(a) Maximum clipped probabilities.
@@ 3.2 The More the Merrier: Dynamic Sampling
Existing RL algorithm suffers from the gradient-decreasing problem when some prompts have accuracy equal to 1. For example for GRPO, if all outputs $\{o_{i}\}_{i=1}^{G}$ of a particular prompt are correct and receive the same reward 1, the resulting advantage for this group is zero. A zero advantage results in no gradients for policy updates, thereby reducing sample efficiency. Empirically, the number of samples with accuracy equal to 1 continues to increase, as shown in Figure 3(b). This means that the effective number of prompts in each batch keeps decreasing, which can lead to larger variance in gradient and dampens the gradient signals for model training.
To this end, we propose to over-sample and filter out prompts with the accuracy equal to 1 and 0 illustrated in Equation 11, leaving all prompts in the batch with effective gradients and keeping a consistent number of prompts. Before training, we keep sampling until the batch is fully filled with samples whose accuracy is neither 0 nor 1.
@@ 3.5 Dataset Transformation
Our dataset is sourced from the AoPS [^2] https://artofproblemsolving.com/ website and official competition homepages through a combination of web scraping and manual annotation. The answers of math dataset typically come in a variety of formats, such as expression, formula and number, which makes it challenging to design comprehensive rules to parse them. To provide accurate reward signals using rules and minimize errors introduced by formula parsers, inspired by AIME, we select and transform the answers into integers, which are easy to parse. For example, if the original answer is expressed in the form of $\frac{a+\sqrt{b}}{c}$, we instruct the LLM to modify the question so that the expected answer becomes $a+b+c$. After selection and transformation, we obtained the DAPO-Math-17K dataset, which consists of 17K prompts, each paired with an integer as the answer.
@@ 5 Conclusion
Tiantian Fan1, Gaohong Liu1, Lingjun Liu1, Xin Liu1, Haibin Lin1, Zhiqi Lin1, Bole Ma1, Guangming Sheng1,3, Yuxuan Tong1,2,4, Qiying Yu1,2,4, Chi Zhang1, Mofan Zhang1, Wang Zhang1, Hang Zhu1, Jinhua Zhu1
@@
Jiaze Chen1, Jiangjie Chen1,4, Chengyi Wang1, Hongli Yu1,2,4, Weinan Dai1,2,4, Yuxuan Song1,2,4, Xiangpeng Wei1, Qiying Yu1,2,4
@@ Acknowledgments
We thank Zhengyin Du, Kai Shen, Tianyang Zhan, Zhen Xiao, Renjie Zheng, Li Han, Kaihua Jiang as well as other colleagues at ByteDance for their support for the DAPO project.

### v2

@@ Abstract
Figure 1: AIME 2024 scores of DAPO on the Qwen2.5-32B base model, outperforming the previous SoTA DeepSeek-R1-Zero-Qwen-32B using 50% training steps. The x-axis represents the gradient update steps.
@@ 3.1 Raise the Ceiling: Clip-Higher
We propose the Clip-Higher strategy to address this issue. Clipping over the importance sampling ratio is introduced in Clipped Proximal Policy Optimization (PPO-Clip) [21] to restrict the trust region and enhance the stability of RL. We identify that the upper clip can restrict the exploration of the policy, where making an ‘exploitation’ token more probable is much easier yet the probability of an unlikely ‘exploration’ token is too tightly bounded to be uplifted.
Concretely, when $\varepsilon=0.2$ (the default value of most algorithms) and $\hat{A}_{i,t}>0$ (the system tries to increase the probability), consider two actions with probabilities $\pi_{\theta_{\text{old}}}(o_{i}\mid q)=0.01$ and $0.9$. The upper bounds of the increased probabilities $\pi_{\theta}(o_{i}\mid q)$ are $0.012$ and $1.08$, respectively $\left(\pi_{\theta_{\text{old}}}\cdot(1+\epsilon)\right)$. This implies that ‘exploitation’ tokens with a higher probability (e.g., 0.9) are not constrained to get even extremely larger probabilities like 0.999. Conversely, for low-probability ‘exploration’ tokens, achieving a non-trivial increase in probability is considerably more challenging. Empirically, we also observe that the mean probability of up-clipped tokens is low: $\pi_{\theta}(o_{i}\mid q)<0.2$ (Figure 3(a)). This finding supports our intuition that the upper clipping threshold indeed restricts the probability increase of low-probability ‘exploration’ tokens, thereby potentially constraining the exploration of the system.
@@
We increase the value of $\varepsilon_{\text{high}}$ to leave more room for the increase of low-probability tokens. As shown in Figure 2, this adjustment effectively enhances the policy’s entropy and facilitates the generation of more diverse samples. We keep $\varepsilon_{\text{low}}$ as it is, because increasing it will suppress the probability of these tokens to 00, resulting in the collapse of the sampling space.
(a) Mean up-clipped probability.
@@
Figure 3: The mean up-clipped probability as well as the ratio of prompts with accuracy=1.
@@ 3.2 The More the Merrier: Dynamic Sampling
Existing RL algorithm suffers from the gradient-decreasing problem when some prompts have accuracy equal to 1. For example for GRPO, if all outputs $\{o_{i}\}_{i=1}^{G}$ of a particular prompt are correct and receive the same reward, the resulting advantage for this group is zero. A zero advantage results in zero policy gradients, shrinking the magnitude and increasing the noise sensitivity of the batch gradient, thereby degrading sample efficiency. Empirically, the number of samples with accuracy equal to 1 continues to increase, as shown in Figure 3(b). This means that the effective number of prompts in each batch keeps decreasing, which can lead to larger variance in gradient and dampens the gradient signals for model training.
To this end, we propose to over-sample and filter out prompts with the accuracy equal to 1 and 0 as illustrated in Equation 11, leaving all prompts in the batch with effective gradients and keeping a consistent number of prompts. The sampling cost for each batch is dynamic. Before training, we keep sampling until the batch is fully filled with samples whose accuracy is neither 0 nor 1.
@@ 3.5 Dataset Transformation
Our dataset is sourced from the web and official competition homepages through a combination of web scraping and manual annotation. The answers of math dataset typically come in a variety of formats, such as expression, formula and number, which makes it challenging to design comprehensive rules to parse them. To provide accurate reward signals using rules and minimize errors introduced by formula parsers, inspired by AIME, we select and transform the answers into integers, which are easy to parse. For example, if the original answer is expressed in the form of $\frac{a+\sqrt{b}}{c}$, we instruct the LLM to modify the question so that the expected answer becomes $a+b+c$. After selection and transformation, we obtained the DAPO-Math-17K dataset, which consists of 17K prompts, each paired with an integer as the answer.
@@ 5 Conclusion
Weinan Dai1,2,4, Tiantian Fan1, Gaohong Liu1, Juncai Liu1, Lingjun Liu1, Xin Liu1, Haibin Lin1, Zhiqi Lin1, Bole Ma1, Guangming Sheng1,3, Yuxuan Tong1,2,4, Qiying Yu1,2,4, Chi Zhang1, Mofan Zhang1, Ru Zhang1, Wang Zhang1, Hang Zhu1, Jinhua Zhu1
@@
Jiaze Chen1, Jiangjie Chen1,4, Chengyi Wang1, Hongli Yu1,2,4, Yuxuan Song1,2,4, Xiangpeng Wei1, Qiying Yu1,2,4
@@ Acknowledgments
We thank Zhengyin Du, Shengding Hu, Kai Shen, Tianyang Zhan, Zhen Xiao, Renjie Zheng, Li Han, Kaihua Jiang as well as other colleagues at ByteDance for their support for the DAPO project.